# Reverse Question Answering: Can an LLM Write a Question so Hard (or Bad) that it Can’t Answer?

### v1

@@ Abstract
Question answering (QA)—producing correct answers for input questions—is popular, but we test a reverse question answering (RQA) task: given an input answer, generate a question with that answer. Past work tests QA and RQA separately, but we test them jointly, comparing their difficulty, aiding benchmark design, and assessing reasoning consistency. 16 LLMs run QA and RQA with trivia questions/answers, revealing: 1) Versus QA, LLMs are much less accurate in RQA for numerical answers, but slightly more accurate in RQA for textual answers; 2) LLMs often answer their own invalid questions from RQA accurately in QA, so RQA errors are not from knowledge gaps alone; 3) RQA errors correlate with question difficulty and inversely correlate with answer frequencies in the Dolma corpus; and 4) LLMs struggle to give valid multi-hop questions. By finding question and answer types yielding RQA errors, we suggest improvements for LLM RQA reasoning. [^1] https://github.com/nbalepur/QG-vs-QA
@@ 1 Reversing the Question Answering Task
Question answering (QA) is a long-standing task in NLP Green Jr et al. (1961). Given an input question $q$, QA aims to deduce the correct answer $a$ Reiter (1989). More recently, large language models (LLMs) have been used for the inverse task—given an answer $a$, generate a valid question $q$ to which $a$ is the answer—which we call reverse question answering (RQA).We distinguish from question generation Zhang et al. (2021), which uses an input context to ground the answer. RQA is used in downstream tasks like exam generation Biancini et al. (2024) and brainstorming Xu et al. (2024).
@@
Second, chaining RQA and QA forms a consistency check for LLM reasoning Liu et al. (2024a). RQA—inferring just one of many valid questions—is abductive Abe (1998), while QA—inferring an answer from question premises—is deductive Reiter (1989). Thus, by seeing if $\textsc{QA}(\textsc{RQA}(a))\approx a$, i.e., checking if an LLM can answer its own question from RQA (Fig 1), we can assess LLMs’ logical robustness in abduction and deduction (section 3.2). This analysis can also help determine if LLMs can reliably self-verify Pan et al. (2024) in downstream RQA tasks like writing exams Wang et al. (2018) or proposing research questions Schurz (2008).
@@
Figure 1: GPT-4 consistency check in RQA/QA. The LLM fails to produce a valid question with answer 488 (top), but correctly gives the answer 77 for its own question (bottom).
In numerical domains, LLMs are much less accurate in RQA than QA, especially integers (Fig 1); the accuracy difference when LLMs do these tasks exceed 0.80 for Command-R and LLaMA-3 (section 3.1). Interestingly, in textual domains, the trend reverses, so LLMs are not consistently better generators or validators Li et al. (2024a). We then design a consistency check (section 3.2) to see if LLMs can answer their own RQA questions; numerical RQA inaccuracies are not solely due to knowledge gaps, as LLMs often answer their own invalid questions correctly in QA (33% of cases for Claude-Opus). We then analyze questions from RQA (section 3.3, section 3.4) and find RQA errors occur when LLMs give overly-complex, multi-step questions, giving insights into strategies—like complexity bias mitigation in preference data and calibrating models using difficulty scores—to improve LLM reliability for RQA.
@@ 2 Experimental Setup
1) $\textsc{QA}(q)\rightarrow\hat{a}$: Given a question $q$ with a single answer $a$, the LLM gives an answer $\hat{a}$ for $q$. QA succeeds if $a$ matches $\hat{a}$ semantically. This typical QA setup tests deduction, as the model must reason to the correct answer $a$ based on premises in $q$.
2) $\textsc{RQA}(a)\rightarrow\hat{q}$: Given answer $a$, the LLM must give a question $\hat{q}$. RQA succeeds if the correct answer to $\hat{q}$ is $a$ (verified via oracle, section 2.3). RQA tests abduction, as the model must reason toward one of many valid questions with the answer of $a$.
Below, we describe the datasets (section 2.1), models (section 2.2), and metrics (section 2.3) used for RQA and QA.
@@
2.1 Dataset Collection
We use 4 domains of question/answer pairs $(q,a)$ for QA and RQA inputs, based on answer type of $a$ (Table 1). We group them as numerical (Number, Number+Text) or textual (Easy Fact, Hard Fact), providing varied domains to test QA and RQA.
For $a$ of type (1), $q$ is a random, one-step math operation (what is 118+211?). Other types are from QANTA Rodriguez et al. (2019), an expert-curated dataset of multi-sentence trivia QA pairs. For (3)–(4), $a$ is the answer to sampled QANTA questions, with the last sentence as $q$. We use middle school questions for (3) and college questions for (4) to discern between easy/hard facts. We find type (2) $a$ in QANTA via regex and $q$ from the sentence $a$ appears in. We verify all QA pairs (Appendix A.1).
2.2 Models
We test 16 LLMs: GPT (Achiam et al., 2023, 3.5, 4, 4o), Command R (Cohere, 2023, Command-R, Command-R+), Claude (Anthropic, 2023, Sonnet, Haiku, Opus), LLaMA-3 Instruct (Dubey et al., 2024, 8B, 70B), Yi-1.5 Chat (Young et al., 2024, 6B, 9B, 34B), and Mistral Instruct (Jiang et al., 2024, 7B, 8x7B, 8x22B). All use 0 temperature.
The QA and RQA prompts are 0-shot, as few-shot exemplars test inductive reasoning, not deduction/abduction Liu et al. (2024a) in QA/RQA. Exemplars also do not improve LLM RQA accuracy (Appendix A.6). Prompts follow the same template as Figure 1 with format rules to parse outputs Liu et al. (2024c). Two NLP graduate students write the prompts, with all design steps in Appendix A.2.
2.3 Evaluation Metrics
To compute QA accuracy, two graduate students researching QA annotate if 1280 LLM QA answers $\hat{a}$ for a question $q$ match its true answer $a$ (20 per answer type/model). We test seven metrics Li et al. (2024b) that check if $\hat{a}$ and $a$ are equivalent. We select DSPy-optimized Khattab et al. (2024) GPT-4o for easy/hard entities and a rule-based method for numerical entities, since these methods had the highest agreement with humans (94% on average).
For RQA accuracy, students annotate if the answer to 1280 questions $\hat{q}$ from RQA is $a$ (20 per answer type/model), following rules from Li et al. (2024b). We use DSPy-optimized GPT-4o as an oracle ($\textsc{Verify}^{*}(\hat{q},a)$) to verify if $a$ is the answer to $\hat{q}$, which has high ($90\%$) human agreement. Metric agreement is high but imperfect, so we also show QA/RQA accuracy via our 1280 annotations (Figure 6), which has the same trend as our metrics.
@@ 3.1 LLMs Struggle with Numerical RQA
We first see if RQA (red, no stripe) or QA (blue, striped) is consistently harder for LLMs (Figure 2). In numerical domains (Number, Number+Text), LLMs are much more accurate in QA versus RQA, revealing a clear abduction weakness. Interestingly, in text domains (Easy, Hard), the trend reverses—RQA slightly beats QA 31/32 times. Thus, LLMs cannot be categorized as always stronger in generation or validation West et al. (2023); Li et al. (2024a): their abilities are domain-specific. If users (e.g. teachers) want to write question-answer pairs with LLMs, we advise manually writing questions for numerical pairs and answers for text pairs, and using LLMs to generate the counterparts, given their strengths in numerical QA and text RQA.
The Numbers domain has the largest QA/RQA accuracy gaps, over 0.8 for LLaMA and Command-R. Some view LLMs as strong math reasoners, but they excel just in deductive QA tasks, as QA is the main testbed for math abilities Ahn et al. (2024). In contrast, abduction in textual domains appears in instruction-tuning datasets with queries like “Tell me about Germany.” Thus, researchers should design more abductive math benchmarks, like RQA, to holistically evaluate LLM math capabilities.
@@ 3.2 QA Can Self-Verify Numerical RQA
We chain RQA and QA for consistency, i.e., see if $\textsc{QA}(\textsc{AQ}(a))\approx a$ (Fig 1). If the check fails, the RQA question $\hat{q}$ is invalid, the LLM fails to answer its own valid $\hat{q}$, or both fail, disentangled below.
@@
LLMs are fairly consistent in textual domains, but often fail the check in numerical domains, except GPT-4o (Figure 3, left). Thus, our LLMs are logically inconsistent in numerical abduction and deduction. In such cases, QA rarely fails alone: either both RQA and QA fail, where the LLM gives an invalid question that it cannot answer, or just RQA fails, where the LLM detects its error. The latter is akin to hallucination snowballing Zhang et al. (2023)—inaccurate questions are not just due to knowledge gaps, as the model can still answer its invalid question accurately (e.g. 33% of cases for Opus). Thus, self-verification Weng et al. (2023) could be a useful strategy to verify and improve the correctness of outputs in numerical RQA tasks.
Figure 4: Answer token count and question difficulty of when RQA succeeds/fails, averaged over all LLMs.
Figure 5: Analysis of Number RQA errors. RQA often fails when the LLM tries to give a multi-step question.
@@ 3.3 Number+Text RQA Errs on Rare Entities
To see when LLMs fail in numerical RQA (section 3.1), we test two indicators of RQA error. We first see how often $a$ appears in the Dolma pretraining corpus Soldaini et al. (2024) via infini-gram Liu et al. (2024b), a proxy for the size of all valid questions an LLM must abductively reason over in RQA. Next, section 3.2 hints LLMs may give overly-hard questions (QA+AQ fail), so we use a Prometheus LLM Kim et al. (2024) to get a 1-5 difficulty score for $\hat{q}$. We average metrics pivoted by RQA success/failure.
@@ 3.4 LLMs Fail to Write Multi-Step Questions
For qualitative insights into question types $\hat{q}$ from RQA, we analyze 30 $\hat{q}$ when RQA fails/succeeds in strong LLMs with low RQA accuracy (section 3.1): L-70B, GPT-4, and Opus. Due to page limits, we use the Numbers split, as its similar answers yield $\hat{q}$ with similar patterns, and group $\hat{q}$ as: 1) Single-Step: has one math operation; 2) Multi-Step: has 2+ math operations; 3) Fact-Based: tests factual knowledge; and 4) Metric Error: metric misclassification. In Appendix A.8, we analyze more questions $\hat{q}$ in aspects like question novelty, answerability, similarity across models, and memorization.
When RQA fails, $\hat{q}$ is often multi-step (Fig 5)—combining math and facts (how many legs are on a human, cat, & spider?) or adding primes (Fig 1). In contrast, valid $\hat{q}$ are often single-step (what is $19^{2}$?) or factual McCarthy (1959) (how many days is a leap year? for 366). We believe the habit of multi-step RQA is from preference tuning; users favor a complex output even if it is wrong Wen et al. (2024). Thus, curbing complexity bias in alignment, or multi-hop QA decoding methods Zhao et al. (2021), may improve LLMs in multi-step RQA.
@@ 4 Related Work
LLM Reasoning: Prior works study LLM reasoning to improve accuracy Qiao et al. (2023) or explainability Si et al. (2024). More recently, works see if LLMs can execute diverse reasoning types, such as inductive Bowen et al. (2024); Yang et al. (2024), deductive Sanyal et al. (2022); Mondorf and Plank (2024), and abductive Zhao et al. (2023); Balepur et al. (2024b) reasoning. However, we are the first to target abduction via RQA, which differs from typical question generation setups as we have no access to an input context Zhang et al. (2021).
LLM Consistency: LLMs must be consistent to reliably help users Visani et al. (2022), but LLMs are inconsistent under perturbations like prompt format Sclar et al. (2024), entity reversal Berglund et al. (2024), negation Ravichander et al. (2022); Balepur et al. (2024a), and option order Zheng et al. (2024). Recent work finds inconsistencies in LLM generation and verification Li et al. (2024a), which we reproduce in our QA/RQA consistency check. Similarly, Deb et al. (2023) and Yu et al. (2024) compare LLMs in forwards (QA) and backwards (filling question blanks given an answer) reasoning in math. While Deb et al. (2023) claim that backwards reasoning is abductive, we argue it is deductive as there is just one answer; we more aptly test abduction/deduction consistency via RQA/QA.

### v2

@@ Abstract
Question answering (QA)—giving correct answers to questions—is a popular task, but we test reverse question answering (RQA): for an input answer, give a question with that answer. Past work tests QA and RQA separately, but we test them jointly, comparing their difficulty, aiding benchmark design, and checking reasoning consistency. We run 16 llms on QA and RQA with trivia questions/answers, revealing: 1) Versus QA, llms are much less accurate in RQA for numerical answers, but slightly more accurate in RQA for textual answers; 2) llms often answer their own invalid questions from RQA accurately in QA, so RQA errors are not from knowledge gaps alone; 3) RQA errors correlate with question difficulty and inversely correlate with answer frequencies in the Dolma corpus; and 4) llms struggle to provide valid multi-hop questions. By finding question and answer types that lead to RQA errors, we suggest improvements for llm reasoning.  [^1] Code and data available at https://github.com/nbalepur/Reverse-QA
@@ 1 Reversing the Question Answering Task
Question answering (QA) is a long-standing task in nlp Green Jr et al. (1961). For an input question $q$, QA deduces the correct answer $a$ Reiter (1989). More recently, large language models (llms) do the reverse—given an answer $a$, generate a valid question $q$ to which $a$ is the answer—which we call reverse question answering (RQA). [^2] This definition differs from question generation Zhang et al. (2021), which grounds the answer to an input context. RQA thus can be a part of downstream tasks like exam question generation Biancini et al. (2024) or search query reformulation Dang and Croft (2010).
@@
Second, chaining RQA and QA forms a consistency check for llm reasoning Liu et al. (2024a). RQA—inferring just one of many valid questions—is abductive Abe (1998), while QA—inferring an answer from question premises—is deductive Reiter (1989). Thus, by seeing if $\textsc{QA}(\textsc{RQA}(a))\approx a$, i.e., checking if an llm can answer its own question from RQA (Fig 1), we can assess llms’ logical robustness in abduction and deduction (section 3.2). This analysis can also help determine if llms can reliably self-verify Pan et al. (2024) in downstream RQA tasks like writing exams Wang et al. (2018).
@@
Figure 1: RQA/QA consistency check using GPT-4. The LLM fails to give a valid question with answer 488 (top), but correctly gives the answer 77 for its own question (bottom).
In numerical domains, llms are much less accurate in RQA than QA, especially integers (Fig 1); the accuracy difference when llms do these tasks exceed 0.80 for Command-R and LLaMA-3 (section 3.1). Interestingly, in textual domains, the trend reverses, so llms are not consistently better generators or validators Li et al. (2024a). We then design a consistency check (section 3.2) to see if llms can answer their own RQA questions; numerical RQA failures are not solely due to knowledge gaps, as llms often answer their own invalid questions correctly in QA (33% of cases for Claude-Opus). We then study questions from RQA (section 3.3, section 3.4) and find errors occur when llms give overly-complex, multi-step questions, giving insights into strategies—like complexity bias mitigation in preference data and calibrating models using difficulty scores—to improve llm RQA reliability. Our contributions are:
  1. 1.
    We use Reverse Question Answering (RQA) to test if llms can provide accurate questions for input answers using abductive reasoning.
  2. 2.
    We reveal many llms have a surprising weakness in RQA on numerical entities, struggling on input answers with lower pretraining token counts and when creating multi-hop questions.
  3. 3.
    We design a consistency check between RQA and QA, showing llms answer their own invalid questions from RQA correctly via QA.
@@ 2 Experimental Setup
  • $\textsc{QA}(q)\rightarrow\hat{a}$: Given a question $q$ with a single answer $a$, the llm produces an answer $\hat{a}$ for $q$. QA succeeds if $a$ matches $\hat{a}$ semantically. For example, given the input “What is the name of the polygon with three sides?” for $q$, an llm using QA should give an $\hat{a}$ that matches “triangle” for $a$. This typical QA setup tests deduction, since the model must reason to the correct answer of $a$ based on the premises in $q$.
  • $\textsc{RQA}(a)\rightarrow\hat{q}$: Given an input answer $a$, the llm must produce a question $\hat{q}$. RQA succeeds if the correct answer to $\hat{q}$ is $a$ (verified via oracle, section 2.3). For example, given the input “triangle” for $a$, an llm using RQA could succeed with $\hat{q}$ as “In eight-ball pool, what shape is used to rack the balls?”. RQA tests abduction, as the model must reason toward one of the many valid questions with the answer $a$.
This section describes the datasets (section 2.1), models (section 2.2), and metrics (section 2.3) used for RQA and QA.
@@
2.1 Dataset Collection
We study question/answer pairs $(q,a)$ in four domains for QA and RQA inputs, based on $a$’s answer type (Table 1). We group them as numerical (Number, Number+Text) or textual (Easy Fact, Hard Fact), providing varied domains for testing.
When $a$ is a Number, $q$ is a random, one-step math operation (what is 118+211?). Other types are from qanta Rodriguez et al. (2019), an expert-curated dataset of multi-sentence trivia qa pairs. For Easy and Hard Facts, $a$ is the answer to sampled qanta questions, with the last sentence333qanta questions are paragraph-long and describe a single answer. Sentences in paragraphs are ordered in decreasing difficulty, so we use the last one, forming the easiest question. as $q$. We use middle school questions for Easy Facts and college questions for Hard Facts. We obtain Number+Text answers $a$ in qanta by finding numbers in full questions via regex and $q$ from the sentence $a$ appears in. One PhD student checks all QA pairs to ensure they are accurate (details in Appendix A.1).
2.2 Models
We evaluate 16 llms: GPT (Achiam et al., 2023, 3.5, 4, 4o), Command R (Cohere, 2023, Command-R, Command-R+), Claude (Anthropic, 2023, Sonnet, Haiku, Opus), LLaMA-3 Instruct (Dubey et al., 2024, 8B, 70B), Yi-1.5 Chat (Young et al., 2024, 6B, 9B, 34B), and Mistral Instruct (Jiang et al., 2024, 7B, 8x7B, 8x22B). All llms use temperature 0. We list all parameters in Appendix A.3.
The QA and RQA prompts are zero-shot, since few-shot exemplars test inductive reasoning, not deduction/abduction Liu et al. (2024a) in QA/RQA. Exemplars also do not improve llm RQA accuracy (Appendix A.6). Prompts follow the same template as Figure 1 with format rules to parse outputs Liu et al. (2024c). Two nlp graduate students write the prompts, with all design steps in Appendix A.2.
2.3 Evaluation Metrics
To compute QA accuracy, two graduate students annotate if 1280 llm QA answers $\hat{a}$ for a question $q$ match its true answer $a$ (20 per answer type/model). [^4] The 1280 total annotations are derived from 16 llms, 4 splits, and 20 annotations on each llm/split combination. We test seven metrics Li et al. (2024b) that evaluate if $\hat{a}$ and $a$ are equivalent. We select DSPy-optimized Khattab et al. (2024) GPT-4o for easy/hard entities and a rule-based method for numerical entities, since these methods had the highest agreement with humans (94% on average).
For RQA accuracy, students annotate if the answer to 1280 questions $\hat{q}$ from RQA is $a$ (20 per answer type/model), following rules from Li et al. (2024b). We use DSPy-optimized GPT-4o as an oracle ($\textsc{Verify}^{*}(\hat{q},a)$) to assess if $a$ answers $\hat{q}$, which has high ($90\%$) human agreement. Metric agreement is high but imperfect, so we also show QA/RQA accuracy using our 1280 annotations in Figure 6, which has the same trend as our metrics.
@@ 3 Evaluation of QA and RQA
Having designed our tasks (section 2), this section tests llms abilities in QA and RQA. llms struggle in RQA on numerical entities (section 3.1) but surprisingly can often detect their own errors (section 3.2). We study the types of entities that lead to RQA errors (section 3.3) and qualitatively analyze differences between accurate and inaccurate questions from RQA (section 3.4).
@@
We first see if RQA (red, no stripe) or QA (blue, striped) is consistently harder for llms (Figure 2). In numerical domains (Number, Number+Text), llms are much more accurate in QA versus RQA, revealing a clear abduction weakness. Interestingly, in text domains (Easy, Hard), the trend reverses—RQA slightly beats QA 31/32 times. Thus, llms cannot be categorized as always stronger in generation or validation West et al. (2023); Li et al. (2024a): their abilities are domain-specific. If users (e.g. teachers) want to write question-answer pairs with llms, we advise manually writing questions for numerical pairs and answers for text pairs, and using llms to generate the counterparts, given their strengths in numerical QA and textual RQA.
The Numbers domain has the largest QA/RQA accuracy gaps, over 0.8 for LLaMA and Command-R. Some view llms as strong math reasoners, but they excel just in deductive QA tasks, as QA is the main testbed for math abilities Ahn et al. (2024). In contrast, abduction in textual domains appears in instruction-tuning datasets with queries like “Tell me about Germany”. Thus, researchers should design more abductive math benchmarks, like RQA, to holistically evaluate llm math capabilities.
@@ 3.2 QA Can Self-Verify Numerical RQA
We chain RQA and QA for consistency, i.e., see if $\textsc{{QA}}(\textsc{RQA}(a))\approx a$ (Figure 1). If the check fails, the RQA question $\hat{q}$ is invalid, the llm fails to answer its own valid $\hat{q}$, or both failures occur. We discuss how we disentangle these cases below.
llms give: 1) a question $\hat{q}$ with answer $a$; and 2) an answer $\hat{a}$ to their own $\hat{q}$ without using $a$. We find three yes/no judgments via our metrics (section 2.3): a) does $a$ answer $\hat{q}$ (RQA succeeds); b) does $\hat{a}$ answer $\hat{q}$ (QA succeeds); and c) are $a$ and $\hat{a}$ equivalent? Answers $\mathcal{A}=(\text{a},\text{b},\text{c})$ to these judgments form a truth table to diagnose llm inconsistencies, which in 91% of cases, fall into the four cases of $\mathcal{A}$ below:
  1. 1.
@@
  2. 2.
[…]
  3. 3.
[…]
  4. 4.
[…]
Other rare cases of $\mathcal{A}$ are metric prediction errors or errors in $\hat{q}$ (e.g. ambiguity), which we omit for this analysis. Appendix A.7 shows all cases of $\mathcal{A}$.
llms are fairly consistent in textual domains, but often fail the check in numerical domains, except GPT-4o (Figure 3, left). Thus, our llms are logically inconsistent in numerical abduction and deduction. In such cases, QA rarely fails alone: either both RQA and QA fail, where the llm gives an invalid question that it cannot answer, or just RQA fails, where the llm detects its error. The latter is akin to hallucination snowballing Zhang et al. (2024)—inaccurate questions are not just due to knowledge gaps, as llm can answer their invalid question accurately (e.g. 33% of cases for Opus).
For instance, given the answer “127 countries”, Opus incorrectly produces the question “How many countries are members of the United Nations that do not have veto power in the UN Security Council?”. However, when Opus answers its own question, it knows there are 193 countries in the UN and five of them have veto power, [^5] At the time of writing this paper. returning the correct answer of “188”. Thus, self-verification Weng et al. (2023) could be a useful way to verify the correctness of responses in numerical RQA tasks.
Figure 4: Answer answer token count in Dolma and question difficulty of when RQA succeeds/fails, averaged over llms.
@@ 3.3 Number+Text RQA Errs on Rare Entities
To find when llms fail in numerical RQA (section 3.1), we test two indicators of RQA error. We first see how often $a$ appears in the Dolma pretraining corpus Soldaini et al. (2024) via infini-gram Liu et al. (2024b), a proxy for the size of all valid questions an llm must abductively reason over in RQA. Next, as section 3.2 hints llms may give overly-hard questions (RQA+QA fail), we use the Prometheus llm Kim et al. (2024) to get a 1-5 difficulty score for $\hat{q}$. We average metrics pivoted by RQA success/failure on the subset containing human annotations (section 2.3).
@@
For qualitative insights into question types $\hat{q}$ from RQA, we analyze 30 $\hat{q}$ when RQA fails/succeeds in strong llms with low RQA accuracy (section 3.1): L-70B, GPT-4, and Opus. For brevity, we just study the Numbers split, as its similar answers yield $\hat{q}$ with similar patterns, and group $\hat{q}$ as: 1) Single-Step: has one math operation; 2) Multi-Step: has 2+ math operations; 3) Fact-Based: tests factual knowledge; and 4) Metric Error: metric misclassification. In Appendix A.8, we analyze more questions $\hat{q}$ in aspects like question novelty, answerability, similarity across models, and memorization.
Figure 5: Analysis of Number RQA errors. RQA often fails when the llm tries to give a complex, multi-step question.
When RQA fails, $\hat{q}$ is often multi-step (Fig 5)—combining math and facts (how many legs are on a human, cat, & spider?) or adding primes (Fig 1). In contrast, valid $\hat{q}$ are often single-step (what is $19^{2}$?) or factual McCarthy (1959) (how many days is a leap year? for 366). We believe the errors in multi-step RQA are from preference tuning; users favor a complex output even if it is wrong Wen et al. (2025). Thus, curbing complexity bias in alignment, or multi-hop QA decoding methods Zhao et al. (2021), may improve llms in multi-step RQA.
@@ 4 Related Work
llm Reasoning: Several works have explored llm reasoning to improve accuracy Qiao et al. (2023) or explainability Si et al. (2024). More recently, works explore if llms can execute diverse reasoning strategies, including inductive Bowen et al. (2024); Yang et al. (2024), deductive Sanyal et al. (2022); Mondorf and Plank (2024), and abductive Zhao et al. (2023); Balepur et al. (2024b) reasoning. However, we are the first to pinpoint abduction abilities via RQA, which differs from traditional question generation setups as we do not have access to an input context Zhang et al. (2021).
llm Consistency: llms must be consistent to reliably help users Visani et al. (2022), but llms are inconsistent under perturbations like prompt format Sclar et al. (2024a), entity reversal Berglund et al. (2024), negation Ravichander et al. (2022); Balepur et al. (2024a), and ordering Zheng et al. (2024). Recent work finds inconsistencies in llm generation and verification in math, QA, style transfer, and coding Li et al. (2024a); Gu et al. (2024), which we reproduce via an RQA/QA consistency check. Deb et al. (2023) and Yu et al. (2024) similarly compare llms in forwards (QA) and backwards (filling question blanks for an answer) reasoning in math. While Deb et al. (2023) claim backwards reasoning is abductive, we argue it is deductive as there is just one answer; we more aptly test abduction/deduction consistency via RQA/QA.